# Stochastic modeling of deterministic laser chaos using generator extended dynamic mode decomposition

### v1

@@ I Introduction
Coupled lasers have been used as test beds to observe various synchronization phenomena (for example, see [1, 2, 3, 4, 5, 6, 7, 8, 9] and the review in [10].) An example of synchronization is the leader-laggard relationship. In the leader-laggard relationship, one of two lasers oscillates in advance of the other by a propagation delay time. The laser oscillating in advance is called the leader, and the other laser that oscillates with a delay and follows the leader is called the laggard. Kanno et al. reported the spontaneous exchange of the leader-laggard relationship [11]; there are autonomous and irregular exchanges between the leader and the laggard.
@@
In the present paper, we aim to construct stochastic models from data generated by two mutually coupled lasers; the coupling has a time-delay, which is crucial to cause the synchronization chaos. Figure 1 shows the problem settings of the present paper. As denoted above, the coarse-graining nature in deriving the leader-laggard relationship should guarantee the validity of the probabilistic description, and we must check it. However, as denoted later, it is difficult to construct a stochastic model describing the whole dynamics of the leader-laggard relationship because the dynamics are quite complicated. Hence, one should employ appropriate data preprocessing, and the preprocessing contributes to clarifying the essential components of the leader-laggard relationship. To construct the stochastic models from the original data, we employ a method in [32], which is based on the Koopman operator approach. The Koopman operator approach enables us to deal with nonlinear systems in terms of linear algebra; we will briefly review the related works in Sec. III. As for the characteristic statistics for comparison, we focus on two distributions, i.e., the distribution of being the leader and that of switching-time intervals, because they could be essential to achieve reinforcement learning in the MAB problem. After discussing the data preprocessing and features of constructed stochastic models, we finally yield a demonstration for the MAB problem using one of the constructed stochastic models.
@@ II.1 Time-delayed mutually coupled semiconductor lasers
$\quad+\kappa_{n+1}E_{n+1}(t-\tau)e^{i\theta_{n}(t)},$ (1)
@@
where $n\in\{1,2\}$ corresponds to the $n$-th laser, and $n=3$ means the first laser, i.e., $E_{3}(t)\equiv E_{1}(t)$ and $N_{3}(t)\equiv N_{1}(t)$. The symbol $\theta_{n}(t)$ is the optical phase difference between the laser light and the injected light, and $\omega_{1}=2\pi c/\lambda_{1}$ is the optical angular frequency for the first laser. Note that $\Delta\omega=2\pi\Delta f_{\textrm{ini}}$ and $\omega_{2}=\omega_{1}-\Delta\omega$. The symbol $\Delta f_{\textrm{ini}}$ represents the initial optical frequency detuning between the two lasers; $\Delta f_{\textrm{ini}}$ is an important variable to control the characteristics of the leader-laggard relationship. The coupling strengths $\kappa_{1}$ and $\kappa_{2}$ are also crucial, and we will change them in the numerical experiments described later. The meanings of other symbols and their values are summarized in Table 1. Here, we basically employ the same settings with [11].
Figure 2: Sample trajectories of optical intensities for the first laser (a) and second laser (b). The vertical axis is drawn using arbitrary units, and the unit of the horizontal axis is $[\mathrm{ns}]$. The initial optical frequency detuning $\Delta f_{\mathrm{ini}}$ is set to $1$ GHz.
@@
Figure 2 shows examples of the intensities of the $n$-th laser, $I_{n}(t)$, where $I_{n}(t)=\lvert E_{n}(t)\rvert^{2}$. Since plots with $\Delta t=1.0\times 10^{-12}\,\mathrm{s}$ are too dense, we plotted the trajectories with the time-interval $1.0\times 10^{-10}\,\mathrm{s}$. Note that the behavior of the first laser (Fig. 2(a)) around $t=25$ is similar to that of the second laser (Fig. 2(b)) around $61(\simeq 25+\tau)$. Hence, there would be a synchronization of the laser chaos. Thus, we see that the intensity of the second laser would lag behind that of the first laser by the propagation delay time $\tau$.
@@ II.2 Cross-correlation and leader-laggard relationship
For applications in machine learning of coupled lasers, short-term correlations are used rather than original trajectories in Fig. 2. The short-term cross-correlation values for the first laser, $C_{1}(t)$, and those for the second laser, $C_{2}(t)$, are given as follows [11]:
@@
The comparison of the cross-correlation values $C_{1}$ and $C_{2}$ leads to the leader-laggard relationship. That is, when $C_{1}>C_{2}$, one considers that $I_{1}(t-\tau)$ and the time-delayed value $I_{2}(t)$ show large synchronization. This indicates that the first laser is the leader. Conversely, $C_{2}>C_{1}$ means that the second laser is the leader. Hence, comparing $C_{1}(t)$ with $C_{2}(t)$ immediately helps us to determine the leader-laggard relationship.
@@
  • There are periodic switching of the leader-laggard relationship with the period $\tau$.
@@ III.2 Brief summary of gEDMD
$\widetilde{\bm{b}}(\bm{x}_{\ell})$ $=\frac{1}{Z_{\bm{x}_{\ell}}}\sum_{n\in\mathcal{C}_{p}}K_{\bm{H}}(\bm{x}_{\ell},\bm{x}_{n})\left[\frac{1}{\Delta t}(\bm{x}_{n+1}-\bm{x}_{n})\right],$ (12)
$\widetilde{A}(\bm{x}_{\ell})$ $=\frac{1}{Z_{\bm{x}}}\sum_{n\in\mathcal{C}_{p}}K_{\bm{H}}(\bm{x}_{\ell},\bm{x}_{n})\left[\frac{1}{\Delta t}(\bm{x}_{n+1}-\bm{x}_{n})(\bm{x}_{n+1}-\bm{x}_{n})^{\top}\right],$ (13)
$Z_{\bm{x}_{\ell}}$ $=\sum_{n\in\mathcal{C}_{p}}K_{\bm{H}}(\bm{x}_{\ell},\bm{x}_{n}),$ (14)
@@ IV.2 Filtered trajectories
  • The preprocessing “filtering with the low-pass filter” focuses on reproducing the probability of being the leader by filtering out the high-frequency comments of the data. This approach is intended to capture the dynamics associated with the probability of being the leader.
Figures 4(b), 4(c), and 4(d) show the preprocessed trajectories. After the preprocessing, we construct stochastic models for the preprocessed data in Figs 4(b), 4(c), and 4(d), respectively. For the LASSO algorithm, we used the hyperparameter $\lambda=1.0\times 10^{-7}$, $3.0\times 10^{-6}$, and $1.0\times 10^{-7}$, for the trajectories in Figs 4(b), 4(c), and 4(d), respectively. Note that we tried several hyperparameter values for each case and selected parameters to yield reasonable stochastic models.
@@ IV.3 Features of statistics
Figure 5: The distribution of being the leader. The distribution evaluated by the original trajectory is (a). By contrast, (b), (c), and (d) are evaluated by the artificial dataset generated by the stochastic models constructed from the trajectories in Figs. 4(b), 4(c), and 4(d), respectively. The vertical gray lines correspond to $C_{1}-C_{2}=0.0$.
@@
After the construction of the stochastic models, we generate artificial datasets using the constructed models. Monte Carlo simulations with the Euler-Maruyama scheme are used for the time-evolution of the stochastic differential equations; the time-interval for the Euler-Maruyama scheme is $\Delta t=0.1\times 10^{-9}\,\mathrm{s}$. In the Monte Carlo simulations, there were exceptional cases where the coordinates tended to diverge. Hence, we employed the following experimental setting: When the simulated trajectory deviated from the region within the original data, we regenerated the noise term; if the regeneration procedure was repeated 10 times, we restarted the time-evolution with randomly selected initial coordinates. Using these generated data points, we evaluate distributions of the following two statistics:
  • The distribution of being the leader. Here, we evaluate the number of data points with $C_{1}(t)>C_{2}(t)$, i.e., the number of times that the first laser becomes the leader.
@@
The distribution of being the leader is crucial for achieving reinforcement learning for the MAB problem. Figure 5(a) shows the distribution of being the leader for the original trajectory in Fig. 4(a), in which we can see a positive peak shift. In the numerical simulation, the first laser tends to be the leader because we used $\Delta f_{\mathrm{ini}}=1.0\times 10^{9}$ Hz. Note that a negative $\Delta f_{\mathrm{ini}}$ leads to the peak shift toward a negative direction. In addition, larger values of $\Delta f_{\mathrm{ini}}$ cause larger peak shifts. These characteristics have already been reported in [11]. In [26], this feature of the peak shift from control parameters is exploited to solve the reinforcement learning tasks, where $\kappa_{1}$ and $\kappa_{2}$ were used as the control parameters instead of $\Delta f_{\mathrm{ini}}$.
@@
The distributions evaluated by artificial datasets from the stochastic models are shown in Figs. 5(b), 5(c), and 5(d). Note that we see the peak shifts in Figs. 5(b) and 5(d), while there is no peak shift in Fig. 5(c). Hence, the high-pass filtered data cannot preserve this essential characteristic. The peak shift needs a bias on the state space on $C_{1}(t)$ and $C_{2}(t)$; in Figs. 4(a), 4(b), and 4(d), we see a bias of the data toward the lower right on the diagonal line. These facts indicate that high-frequency noise focused on the high-pass filtered data is not crucial in recovering the leader-laggard relationship.
@@
Figure 6 shows the distributions of leader switching-time intervals from Laser 1 to Laser 2. Note that we depict these figures based on the histogram. Hence, even if the values on the horizontal axis are close, there may be only one data point in a bin on the histogram. If so, the value of the probability becomes small. Therefore, we also colored the densities for those with nonzero values using a kernel-density estimation method. The distribution evaluated by the original trajectory, Fig. 6(a), shows a power-law decay and a peak structure away from the power-law. One might consider that the behavior in Fig. 3 is periodic, but there are many switchings with small time intervals, which lead to the power-law decay in Fig. 6(a). Note that we can see the power-law decay in all cases in Figs. 6(b), 6(c) and 6(d). However, the tendency to concentrate apart from the power-law decay is seen only in Figs. 6(a) and 6(d). Although Fig. 6(d) may not look like a peak structure, the yellow area is away from the power-law decay, which corresponds to the peak structure in Fig. 6(a). The peak structure indicates the periodic behavior in Fig. 3. Hence, the trajectories within the high-density cluster region and the high-pass filtered data lose this periodicity. By contrast, the low-pass filtered data remains appropriately preserved for this feature.
@@
Of course, there could be other important statistics. Next, we will confirm that the stochastic model constructed by the low-pass filtered data works well in the reinforcement learning task, i.e., the MAB problem.
@@ IV.4 The multi-armed bandit problem
Table 3: Parameter values used in the numerical experiments.
Symbol Parameter Value
$k$ Step width $4$
$N$ Number of step levels for positive $\Delta\kappa$ $16$
$\frac{2N}{k}+1$ Total number of step levels $9$
$a$ memory parameter $0.999$
$P_{1}$ Hit probability of slot machine $1$ $0.6$
$P_{2}$ Hit probability of slot machine $2$ $0.4$
@@
Since the experimental setting for the reinforcement learning is similar to that in [26], we explain the key points briefly. Figure 7 shows the setting. The first and second lasers are assigned to two slot machines, respectively. When the first (second) laser is the leader, the first (second) slot machine is selected. The selected slot machine yields one of the results, hit or miss, and the result affects the control parameters $\kappa_{1}$ and $\kappa_{2}$, i.e., $\Delta\kappa\equiv\kappa_{1}-\kappa_{2}$. The value of $\Delta\kappa(t)$ is updated as follows:
$\Delta\kappa(t)=\begin{cases}\,\,kN&\left(\lceil TA(t)\rceil>N\right),\\ \,\,kTA(t)&\left(-N<\lceil TA(t)\rceil<N\right),\\ \,\,-kN&\left(\lceil TA(t)\rceil<-N\right),\end{cases}$ (25)
@@
$\overline{P}_{i}=\frac{(\textrm{Number of hits for the $i$-th slot machine})}{(\textrm{Number of selections for the $i$-th slot machine})}.$ (27)
@@
We here perform the computation $1000$ plays per cycle and repeat this for $10$ cycles ($n=10$) to compute the average over the $10$ runs and calculate average total rewards (ATR), which is defined as follows:
@@
As a result, the ATR values were $0.5912$ for the original deterministic laser chaos case and $0.5952$ for the constructed stochastic model; the estimated hit probabilities, $(\overline{p}_{1},\overline{p}_{2})$, were $(0.5978,0.3419)$ for the original deterministic laser chaos case and $(0.6004,0.4517)$ for the constructed stochastic model. Note that the second slot machine is not selected finally, and the estimation $\overline{p}_{2}$ could vary. However, the estimation for the hit probability $\overline{p}_{1}$ works well. Then, we conclude that the constructed stochastic model achieves the same performance as the original deterministic laser chaos in the reinforcement learning problem.

### v3

@@ I Introduction
Coupled lasers have been used as test beds to observe various synchronization phenomena (for example, see [1, 2, 3, 4, 5, 6, 7, 8, 9] and the review in [10].) An example of synchronization is the leader-laggard relationship. In the leader-laggard relationship, one of two lasers oscillates in advance of the other by a propagation delay time. The laser oscillating in advance is called the leader, and the other one that oscillates with a delay and follows the leader is called the laggard. Kanno et al. reported the spontaneous exchange of the leader-laggard relationship [11]; there are autonomous and irregular exchanges between the leader and the laggard.
@@
In the present paper, we aim to construct stochastic models from data generated by two mutually coupled lasers; the coupling has a time-delay, which is crucial to cause the synchronization chaos. Figure 1 shows the problem settings of the present paper. As denoted above, the coarse-graining nature in deriving the leader-laggard relationship should guarantee the validity of the probabilistic description, and we must check it. However, as denoted later, it is difficult to construct a stochastic model describing the whole dynamics of the leader-laggard relationship because the dynamics are quite complicated. The key point here is not to perfectly reproduce the behavior, but to reproduce the features that play a crucial role in reinforcement learning. Hence, one should employ appropriate data preprocessing, and the preprocessing contributes to clarifying the essential components of the leader-laggard relationship in the MAB problem. To construct the stochastic models from data, we employ a method in [32], which is based on the Koopman operator approach. The Koopman operator approach enables us to deal with nonlinear systems in terms of linear algebra; we will briefly review the related works in Sec. III. As for the characteristic statistics for comparison, we focus on two distributions, i.e., the distribution of being the leader and that of switching-time intervals, because they could be essential to achieve reinforcement learning in the MAB problem. After discussing the data preprocessing and features of constructed stochastic models, we finally yield a demonstration for the MAB problem using one of the constructed stochastic models.
@@ II.1 Time-delayed mutually coupled semiconductor lasers
$\qquad+\kappa_{n+1}E_{n+1}(t-\tau)e^{i\theta_{n}(t)},$ (1)
@@
where $n\in\{1,2\}$ corresponds to the $n$-th laser, and $n=3$ means Laser 1, i.e., $E_{3}(t)\equiv E_{1}(t)$ and $N_{3}(t)\equiv N_{1}(t)$. The symbol $\theta_{n}(t)$ is the optical phase difference between the laser light and the injected light, and $\omega_{1}=2\pi c/\lambda_{1}$ is the optical angular frequency for the first laser. Note that $\Delta\omega=2\pi\Delta f_{\textrm{ini}}$ and $\omega_{2}=\omega_{1}-\Delta\omega$. The symbol $\Delta f_{\textrm{ini}}$ represents the initial optical frequency detuning between the two lasers; $\Delta f_{\textrm{ini}}$ is an important variable to control the characteristics of the leader-laggard relationship. The coupling strengths $\kappa_{1}$ and $\kappa_{2}$ are also crucial, and we will change them in the numerical experiments described later. The meanings of other symbols and their values are summarized in Table 1. Here, we basically employ the same settings with [11].
Figure 2: Sample trajectories of optical intensities for Laser 1 (a) and Laser 2 (b). The vertical axis is drawn using arbitrary unit. The initial optical frequency detuning $\Delta f_{\mathrm{ini}}$ is set to 11 GHz.
@@
Figure 2 shows examples of the intensities of the $n$-th laser, $I_{n}(t)$, where $I_{n}(t)=\lvert E_{n}(t)\rvert^{2}$. Since plots with $\Delta t=1.0\times 10^{-12}\,\mathrm{s}$ are too dense, we plotted the trajectories with the time-interval $1.0\times 10^{-10}\,\mathrm{s}$. Note that the behavior of Laser 1 (Fig. 2(a)) around $t=25$ is similar to that of Laser 2 (Fig. 2(b)) around $61(\simeq 25+\tau)$. Hence, there would be a synchronization of the laser chaos. Thus, we see that the intensity of Laser 2 would lag behind that of Laser 1 by the propagation delay time $\tau$.
@@ II.2 Cross-correlation and leader-laggard relationship
For applications in machine learning of coupled lasers, short-term correlations are used rather than original trajectories in Fig. 2. The short-term cross-correlation values for Laser 1, $C_{1}(t)$, and those for Laser 2, $C_{2}(t)$, are given as follows [11]:
@@
The comparison of the cross-correlation values $C_{1}$ and $C_{2}$ leads to the leader-laggard relationship. That is, when $C_{1}>C_{2}$, one considers that $I_{1}(t-\tau)$ and the time-delayed value $I_{2}(t)$ show large synchronization. This indicates that Laser 1 is the leader. Conversely, $C_{2}>C_{1}$ means that Laser 2 is the leader. Hence, comparing $C_{1}(t)$ with $C_{2}(t)$ immediately helps us to determine the leader-laggard relationship.
@@
  • There are periodic switchings of the leader-laggard relationship with the period $\tau$.
@@ III.2 Brief summary of gEDMD
$\widetilde{\bm{b}}(\bm{x}_{\ell})$ $=\frac{1}{Z_{\bm{x}_{\ell}}}\sum_{n\in\mathcal{C}_{p}}K_{\bm{H}_{p}}(\bm{x}_{\ell},\bm{x}_{n})\left[\frac{1}{\Delta t}(\bm{x}_{n+1}-\bm{x}_{n})\right],$ (12)
$\widetilde{A}(\bm{x}_{\ell})$ $=\frac{1}{Z_{\bm{x}}}\sum_{n\in\mathcal{C}_{p}}K_{\bm{H}_{p}}(\bm{x}_{\ell},\bm{x}_{n})\left[\frac{1}{\Delta t}(\bm{x}_{n+1}-\bm{x}_{n})(\bm{x}_{n+1}-\bm{x}_{n})^{\top}\right],$ (13)
$Z_{\bm{x}_{\ell}}$ $=\sum_{n\in\mathcal{C}_{p}}K_{\bm{H}_{p}}(\bm{x}_{\ell},\bm{x}_{n}),$ (14)
@@ IV.2 Filtered trajectories
  • The preprocessing “filtering with the low-pass filter” focuses on reproducing the probability of being the leader by filtering out the high-frequency components of the data. This approach is intended to capture the dynamics associated with the probability of being the leader.
Note that the high-pass and low-pass filters are applied to the cross-correlation trajectory dataset in Fig. 4(a). In [11], the filter was applied to the intensity of the original laser chaos, as in Fig. 2.
Figures 4(b), 4(c), and 4(d) show the preprocessed trajectories. After the preprocessing, we construct stochastic models for the preprocessed data in Figs 4(b), 4(c), and 4(d), respectively. For the LASSO algorithm, we used the hyperparameter $\lambda=3.0\times 10^{-9}$, $3.0\times 10^{-6}$, and $5.0\times 10^{-7}$, for the trajectories in Figs 4(b), 4(c), and 4(d), respectively. Note that we tried several hyperparameter values for each case and selected parameters to yield reasonable stochastic models.
@@ IV.3 Features of statistics
Figure 5: The distribution of being the leader. The distribution evaluated by the original trajectory is (a). By contrast, (b), (c), and (d) are evaluated by the artificial dataset generated by the stochastic models constructed from the trajectories in Figs. 4(b), 4(c), and 4(d), respectively. The blue lines correspond to the histograms, and the orange curves are obtained from the kernel density estimation. The vertical gray lines correspond to $C_{1}-C_{2}=0.0$.
@@
After the construction of the stochastic models, we generate artificial datasets using the constructed models. Monte Carlo simulations with the Euler-Maruyama scheme are used for the time-evolution of the stochastic differential equations; the time-interval for the Euler-Maruyama scheme is $\Delta t=1.0\times 10^{-10}\,\mathrm{s}$. In the Monte Carlo simulations, there were exceptional cases where the coordinates tended to diverge. Hence, we employed the following experimental setting: When the simulated trajectory deviated from the region within the original data, we regenerated the noise term; if the regeneration procedure was repeated 10 times, we restarted the time-evolution with randomly selected initial coordinates. Using these generated data points, we evaluate distributions of the following two statistics:
  • The distribution of being the leader. Here, we evaluate the number of data points with $C_{1}(t)>C_{2}(t)$, i.e., the number of times that Laser 1 becomes the leader.
@@
The distribution of being the leader is crucial for achieving reinforcement learning for the MAB problem. Figure 5(a) shows the distribution of being the leader for the original trajectory in Fig. 4(a), in which we can see a positive peak shift. In the figures, we depict not only the histograms, but also the curves which are obtained from the kernel density estimation with scipy.stats.gaussian_kde; the bandwidth is determined by the Scott’s rule $n^{(-1/(d+4))}$, where $n$ is the number of data points and $d$ is the number of dimensions. In the numerical simulation, Laser 1 tends to be the leader because we used $\Delta f_{\mathrm{ini}}=1.0\times 10^{9}$ Hz. Note that a negative $\Delta f_{\mathrm{ini}}$ leads to the peak shift toward a negative direction. In addition, larger values of $\Delta f_{\mathrm{ini}}$ cause larger peak shifts. These characteristics have already been reported in [11]. In [26], this feature of the peak shift from control parameters is exploited to solve the reinforcement learning tasks, where $\kappa_{1}$ and $\kappa_{2}$ were used as the control parameters instead of $\Delta f_{\mathrm{ini}}$.
@@
The distributions evaluated by artificial datasets from the stochastic models are shown in Figs. 5(b), 5(c), and 5(d). Since we will employ the dependency between the peak in $C_{1}-C_{2}$ and the initial frequency difference $\Delta f_{\mathrm{ini}}$ in the MAB problem, as denoted above, we here focus on the peak shifts in Figs. 5(b) and 5(d). Then, there is no peak shift in Fig. 5(c), and hence, the high-pass filtered data cannot preserve this essential characteristic. The peak shift needs a bias on the state space on $C_{1}(t)$ and $C_{2}(t)$; in Figs. 4(a), 4(b), and 4(d), we see a bias of the data toward the lower right on the diagonal line. These facts indicate that high-frequency noise focused on the high-pass filtered data is not crucial in recovering the leader-laggard relationship. Here, we comment on the complicated shape in Fig. 5(d). The main focus here is not on the complete recovery of statistical properties. Although the preprocessing with the low-pass filter yields a complicated shape, it is crucial to investigate whether the preprocessing preserves the essential features for the MAB problem. We will discuss this point in Sec. IV.D.
Figure 6 shows the distributions of leader switching-time intervals from Laser 1 to Laser 2. Here, the leader switching-time interval means the time-interval at which there are switching from $C_{1}-C_{2}\geq 0$ to $C_{1}-C_{2}<0$. First, we construct the histogram from the simulated data; the bin width is set to $1.0\times 10^{-10}$ s, which corresponds to the time-interval in the simulation. Since we used the log-log plots in Fig. 6, the bin size is too narrow in the right-hand region of the figures; there may be only one data point in a bin on the histogram. This causes the probability values to become too small. Therefore, we also colored densities with non-zero values using kernel density estimation. We here employed scipy.stats.gaussian_kde for the kernel density estimation. The color of each point in the histogram was changed according to the probability density values obtained from kernel density estimation.
@@
The distribution evaluated by the original trajectory, Fig. 6(a), shows a power-law decay and a peak structure away from the power-law. One might consider that the behavior in Fig. 3 is periodic, but there are many switchings with small time intervals, which lead to the power-law decay in Fig. 6(a). Note that we can see the power-law decay in all cases in Figs. 6(b), 6(c) and 6(d). However, the tendency to concentrate apart from the power-law decay is seen only in Figs. 6(a) and 6(d). Although Fig. 6(d) may not look like a peak structure, the yellow area is away from the power-law decay, which corresponds to the peak structure in Fig. 6(a). The peak structure indicates the periodic behavior in Fig. 3. Hence, the trajectories within the high-density cluster region and the high-pass filtered data lose this periodicity. By contrast, the low-pass filtered data remains appropriately preserved for this feature. We show some numerical results for other initial frequency difference $\Delta f_{\mathrm{ini}}$ in Appendix A.
@@
Of course, there could be other important statistics. Furthermore, while the peak shift structure in Fig. 5(d) matches that in Fig. 5(a), it exhibits a more complex shape. The key point here is not to perfectly reproduce the statistics, but to investigate what constitutes the essential characteristics of laser chaos in reinforcement learning tasks, as denoted above. Next, we will confirm whether the stochastic model constructed by the low-pass filtered data works well in the reinforcement learning task, i.e., the MAB problem.
@@ IV.4 The multi-armed bandit problem
Since the experimental setting for the reinforcement learning is similar to that in [26], we explain the key points briefly. Figure 7 shows the setting. The first and second lasers are assigned to two slot machines, respectively. When Laser 1 (or Laser 2) is the leader, the first (or second) slot machine is selected. The selected slot machine yields one of the results, hit or miss, and the result affects the control parameters $\kappa_{1}$ and $\kappa_{2}$, i.e., $\Delta\kappa\equiv\kappa_{1}-\kappa_{2}$. The value of $\Delta\kappa(t)$ is updated as follows:
$\Delta\kappa(t)=\begin{cases}\,\,kN&\left(\textrm{(int)}TA(t)>N\right),\\ \,\,k\textrm{(int)}TA(t)&\left(-N<\textrm{(int)}TA(t)<N\right),\\ \,\,-kN&\left(\textrm{(int)}TA(t)<-N\right),\end{cases}$ (25)
@@
[…]
@@
Table 3: Parameter values used in the numerical experiments.
Symbol Parameter Value
$k$ Step width 44 [ns-1]
NN Number of step levels for positive $\Delta\kappa$ 44
$\frac{2N}{k}+1$ Total number of step levels 99
$a$ Memory parameter $0.999$
$P_{1}$ Hit probability of slot machine 11 $0.6$
$P_{2}$ Hit probability of slot machine 22 $0.4$
$\tau_{\mathrm{SI}}$ Sampling interval 7.0 [ns]
We here perform the computation $1000$ plays per cycle and repeat this for $100$ cycles ($n=100$) to compute the average over the $100$ runs and calculate average total rewards (ATR), which is defined as follows:
@@
As a result, the ATR values were $0.5700$ for the original deterministic laser chaos case and $0.5751$ for the constructed stochastic model; the estimated hit probabilities, $(\overline{p}_{1},\overline{p}_{2})$, were $(0.5999,0.3986)$ for the original deterministic laser chaos case and $(0.5624,0.3920)$ for the constructed stochastic model. Numerical results with different experimental settings are shown in Appendix B.
In addition, we also examine the average convergence time (ACT); the ATR corresponds to the performance after learning, while the ACT relates to the dynamics of the learning process. Here, we define the ACT as follows: we evaluate the time at which learning converges, i.e., $\Delta\kappa(t)=16$ or 16-16, and take the average of these times as the ACT. For the original deterministic laser chaos case, the ACT value is 112.53; in the case with the constructed stochastic model, we have 115.35 as the ACT value. Then, the constructed stochastic model yields a similar ACT value to that of the original laser chaos data.
From these results, we conclude that the constructed stochastic model achieves the same performance as the original deterministic laser chaos in the reinforcement learning problem.